# Supernematic

### v1

@@ III Searching for new quantum number as a tiling problem
We seek the quantum number that is preserved under local resonating dynamics (Fig.1b) and higher order terms) by mapping the problem of allowed particle configurations to the problem of tiling. Tiling problems are a class of combinatorial problems generally considered NP-hard. The rule of the game is to cover a given bounded region with certain geometric shapes, i.e., the tiles, such that the tiles neither overlap nor have gaps in between. The seminal work of Conway and Lagarias [10] shifted the focus from finding a tiling solution to establishing a no-go theorem of tiling impossibility: tiling is impossible when boundary invariants associated with the tile set are violated, where the boundary invariants are derived using the language of combinatorial group theory. As we will make precise in the following, such a boundary invariant can lead to a new quantum number when dynamics is restricted to configurations satisfying cluster-charging contraints at filling $2/3$.
@@
Now, to further connect to the original problem considered in Ref.[10], we view the bigger tiles with the green/magenta contours in Fig.2 as consisting of the smaller three-in-line tiles and triangular tiles of three cojoining plaquettes (blue shades in 2c). All the allowed particle configurations can therefore be mapped to tiling patterns by connecting the hexagonal pieces in Fig.2c) in a boundary-matching way. The sites with particles are mapped to magenta tiles. Note that all the tiles fall into the tile sets in Fig.3a) and b).
For these tiles, the following theorem proved in Ref.[10] connects allowed tiling configurations to a boundary invariant.
@@ III.1 Conway-Lagarias Invariant
To construct the boundary invariant in Thm.1, Ref.[10] maps the boundary of the tiling to another “space”, where the boundaries of the different tiling types carry different winding numbers.
@@ IV.1 Microscopic mapping between particle occupation and dimers
The cluster-charging constraints can be recast into the more familiar constraints in dimer models by viewing the particle as three conjoined dimers. We used this mapping in Ref.[13] to derive the 2D statistical field theory for the classical cluster-charging model at filling $1/3$, where only two height fields are independent at long wavelength since the fluctuations around the constraint relating the three height fields are irrelevant in the classical theory. We now generalize this mapping to filling $2/3$ and consider the effect of quantum fluctuations. Quantum theory lives in 2+12+1D, and we will find three independent height fields because the quantum fluctuation of the constraint is gapless.
@@
First, we view the dual lattice of the physical honeycomb lattice as a set of three interpenetrating honeycomb lattices, labelled as red, green, and blue in Fig.5 a). We then map the particle occupation to three dimer occupations (red, blue and green dimers in Fig.5 a)). The mapping between the dimer occupation and physical particle number can be written as $d_{a}^{r}=n_{i}+n_{j}$, where $d_{a}^{r}$ denotes the “red” dimer occupation number at bond $a$ and $n_{i,j}$ are the particle numbers at site $i,j$, which are the sites linked by the edge of the physical lattice intersected with bond $a$, as noted in Fig.5 b). The analogous expressions also hold for the green and blue lattices. At filling $2/3$, the cluster-charging constraints impose “two-dimers per site” conditions on the colores honeycomb lattice with double-occupation of dimers allowed (illustrated in Fig.5 b) for the red lattice). The corresponding dimer model of each colored honeycomb lattice is the so-called “double-dimer” model, because the configuration can be thought of as overlaying two dimer coverings with one dimer touching each vertex [20].
@@
Now we introduce the height fields. In the dimer model, height fields are integer-valued fields defined at the centers of the plaquettes to encode the dimer constraints [21, 22]. Since the double-dimer configuration can be decomposed into two single-dimer ones, Ref. [23] introduces two height fields for each of the replicas. Here, for the cluster-charging model, only the sum of the two heights is relevant since no physical operators can distinguish the two replicas. We will call this summation “the height field” in the following. The height fields of the red double-dimer model are defined at the centers of the hexagonal plaquettes, with the following rules (Fig.5 c)): when going around the vertex emitting “Y” shaped edges clockwise, if no dimer/one dimer/two dimers is crossed, height is changed by 2-2/+1+1/+4+4. The height assignment is the same for going around another type of sublattice counterclockwise. Under such rules, the action of the $\mathcal{O}_{2}$ operator is to shift the height field of one color by 3-3 and the height field of another color by +3+3 (Fig.5 d)). Now the total $\tilde{N}\equiv N_{A}-N_{B}$ conservation proved in sectionIII maps to the global conservation of the total height fields $h_{r}+h_{g}+h_{b}$.
@@
where the $\triangleright/\triangleleft$ is the shorthand notation for summing the three height fields surrounding the site of the corresponding particle (see Fig. 5 e)), where the color labels are implicit. The constant $\frac{1}{3}$ takes care of the filling fraction. The representation of the particle number using the fields on dual lattice can also be viewed as a plaquette duality, utilized for studying various spin models with ring exchange interactions [24, 25]. The additional link to the height fields is originated from the cluster-charging constraints. Now, the on-site Hubbard term $H_{0}$ can be written as,
@@ IV.2 Coarse-grained height fields and their properties under lattice symmetries
First, the compactification radius of $h^{r,g,b}$ is $3$ for our double-dimer model, since the uniform shift by three gives the same dimer configuration (same as the dimer model on honeycomb lattice [26]). Therefore, the physical operators can only be $\partial h^{r,g,b}$ (and higher order derivatives), the vertex operators $\mathcal{V}_{\vec{m}}(x)=e^{i\frac{2\pi}{3}\vec{m}\cdot\vec{h}(x)}$, where $\vec{m}=(m^{r},m^{g},m^{b})$ is a vector of integers and $\vec{h}=(h^{r},h^{g},h^{b})$, and their combinations.
The transformation of the coarse-grained fields under lattice symmetries can be derived from the “flat” configurations of $h^{r,g,b}$’s. These are columnar orders of the double-dimer model. However, not all columnar order combinations are physical, since not all double-dimer configurations can be mapped to physical ones. There are nine physical columnar orders in total, related to the configuration in Fig.1 b) by lattice translations and rotations. The symmetry transformation of the coarse-grained height fields need to account for not only the symmetry properties of the microscopic heights, but also the transformation of the average height, analogous to the discussion of the dimer model [26, 27]. The detail is discussed in the Appendix Sec.A. We will quote the results below. Under lattice symmetry,
@@ IV.3 Effective field theory description
$\mathcal{S}_{sg}[\eta]=\int\frac{1}{4\tilde{t}^{\prime 2}}(\partial_{\tau}\vec{\eta})^{2}+\tilde{v}^{\prime}(\partial_{\mu}\vec{\eta})^{2}-\tilde{y}^{\prime}\sum_{i}\cos(4\pi\alpha_{i}\cdot\vec{\eta}).$ (21)
@@
Now let us focus on $\mathcal{L}_{1}$, with $\tilde{y}_{0}\rightarrow\tilde{y}_{0}+\tilde{y}$. This theory can be mapped to the effective description of the 2+12+1D quantum XY model as follows. First, we apply the Villain approximation[28], which approximates the cosine term with a sum of Gaussian functions centered around the integers and then integrates out the $h_{0}$ field. It is readily seen that
@@
where $\tilde{\mathbf{y}}$ is a monotonic function of $\tilde{y}_{0}+\tilde{y}$. The Euclidean action $\mathcal{S}_{1}[\theta_{0}]$ also describes the 3D classical XY model with $\theta_{0}$ as the spin azimuthal angle. By taking into account the vortices of $\theta_{0}$, there is a phase transtion between a superfluid phase and a gapped Mott insulating phases by tuning the effective stiffness $K_{\text{eff}}=\sqrt{\tilde{t}/\tilde{\mathbf{y}}}$ across a critical value $K_{\text{eff}}^{c}$[29]. Hence, there are two distinct phases:
  • For $\tilde{t}<\tilde{\mathbf{y}}(K_{\text{eff}}^{c})^{2}$, $\langle h_{0}\rangle\neq 0,~\langle\eta\rangle\neq 0$: the gapped columnar ordered phase.
  • For $\tilde{t}>\tilde{\mathbf{y}}(K_{\text{eff}}^{c})^{2}$, $\langle\theta_{0}\rangle\neq 0,~\langle\eta\rangle\neq 0$: the gapless supernematic phase.
We discuss the properties of each phase in the following sections.
@@ V Columnar ordered phase
In the classical limit of $\tilde{t}=0$, we reproduce the trimer liquid[13], a classical ensemble of extensively-many degenerate configurations (1 e)). In general, small quantum fluctuations around extensive degeneracy can result in quantum order-by-disorder, where a charge ordering state is favored due to the virtual exchange processes. To determine the symmetry-breaking pattern of such a charge-ordered state and its thermal behavior, we use the effective field theory and the symmetry representations of the fields.
@@ V.2 Finite temperature
(a) Upon increasing temperature, $h_{0}$ melts first, via a three-state Potts transition. The intermediate phase has $\vec{\eta}$ ordering. Since $h_{0}$ is disordered, the order parameters are the lowest order vertex operators involving only $\vec{\eta}$-field: $\langle\mathcal{V}_{(1,-1,0)}\rangle$, $\langle\mathcal{V}_{(0,1,-1)}\rangle$ and $\langle\mathcal{V}_{(-1,0,1)}\rangle$. They all take the same value of $1,e^{i\frac{2\pi}{3}},\text{or}e^{-i\frac{2\pi}{3}}$ from Eq.23. From Eq.18, the ordering pattern breaks $\mathcal{C}_{3}$ rotation while preserving translation and inversion. Hence, this intermediate phase is three-fold degenerate, which we call “nematic”. In the disordered phase of $h_{0}$, the bionic monomers can acquire non-zero fugacity. The scaling dimension of these monomer operators can be obtained from the same method as in the classical trimer liquid [13] and $\Delta_{m_{\eta}}=\frac{3K^{\prime}}{4\pi}$. Hence when $K^{\prime}=K_{c}^{{}^{\prime}}=\frac{8\pi}{3}$, the monomers start to become relevant. However, since $\Delta_{(2,-1,-1)}(K_{c}^{{}^{\prime}})=\frac{3}{2}<2$, the thermal phase transition is dominant by the cosine terms, where the value of the critical $K^{\prime}$ is nonuniversal, and expected to be three-state Potts like [31].
@@ VI.1 Zero temperature
where we ignore the dependence on the gapped $\eta$-fields and a spatially oscillating term (see SM A.3 for the derivation). Following the effective field theory of $\theta_{0}$ (Eq.22) and the standard derivation following the hydrodynamic theory of bosonic superfluid [32] (see SM B), the dynamical structure factor $\tilde{S}(q,\omega)\sim|q|\delta(\omega-v|q|)$, where $\tilde{S}(q,\omega)=\int dte^{i\omega t}\langle\tilde{n}(q,t)\tilde{n}(-q,0)\rangle$, and $\tilde{n}(q)$ is the Fourier transform of $n_{A}(x_{A})-n_{B}(x_{B})$. Therefore, the SN is compressible with respect to the sublattice particle number $\tilde{N}$, following the compressibility sum-rule [33] $\tilde{\kappa}=2\lim_{q\rightarrow 0}\int_{0}^{\infty}\frac{\tilde{S}(q,\omega)}{\omega}d\omega$. Hence, from the $\tilde{N}$ perspective, SN is compressible although it is an incompressible state for the total particle number $N=N_{A}+N_{B}$.

### v2

@@ III Searching for new quantum number as a tiling problem
We seek the quantum number that is preserved under local resonating dynamics (Fig.1b) and higher order terms) by mapping the problem of allowed particle configurations to the problem of tiling. Tiling problems are a class of combinatorial problems generally considered NP-hard. The rule of the game is to cover a given bounded region with certain geometric shapes, i.e., the tiles, such that the tiles neither overlap nor have gaps in between. The seminal work of Conway and Lagarias [conway1990tiling] shifted the focus from finding a tiling solution to establishing a no-go theorem of tiling impossibility: tiling is impossible when boundary invariants associated with the tile set are violated, where the boundary invariants are derived using the language of combinatorial group theory. As we will make precise in the following, such a boundary invariant can lead to a new quantum number when dynamics is restricted to configurations satisfying cluster-charging contraints at filling $2/3$.
@@
Now, to further connect to the original problem considered in Ref.[conway1990tiling], we view the bigger tiles with the green/magenta contours in Fig.2 as consisting of the smaller three-in-line tiles and triangular tiles of three cojoining plaquettes (blue shades in 2c). All the allowed particle configurations can therefore be mapped to tiling patterns by connecting the hexagonal pieces in Fig.2c) in a boundary-matching way. The sites with particles are mapped to magenta tiles. Note that all the tiles fall into the tile sets in Fig.3a) and b).
For these tiles, the following theorem proved in Ref.[conway1990tiling] connects allowed tiling configurations to a boundary invariant.
@@ III.1 Conway-Lagarias Invariant
To construct the boundary invariant in Thm.1, Ref.[conway1990tiling] maps the boundary of the tiling to another “space”, where the boundaries of the different tiling types carry different winding numbers.
@@ IV.1 Microscopic mapping between particle occupation and dimers
The cluster-charging constraints can be recast into the more familiar constraints in dimer models by viewing the particle as three conjoined dimers. We used this mapping in Ref.[zhang2024bionicfractionalizationtrimermodel] to derive the 2D statistical field theory for the classical cluster-charging model at filling $1/3$, where only two height fields are independent at long wavelength since the fluctuations around the constraint relating the three height fields are irrelevant in the classical theory. We now generalize this mapping to filling $2/3$ and consider the effect of quantum fluctuations. Quantum theory lives in 2+12+1D, and we will find three independent height fields because the quantum fluctuation of the constraint is gapless.
@@
First, we view the dual lattice of the physical honeycomb lattice as a set of three interpenetrating honeycomb lattices, labelled as red, green, and blue in Fig.5 a). We then map the particle occupation to three dimer occupations (red, blue and green dimers in Fig.5 a)). The mapping between the dimer occupation and physical particle number can be written as $d_{a}^{r}=n_{i}+n_{j}$, where $d_{a}^{r}$ denotes the “red” dimer occupation number at bond $a$ and $n_{i,j}$ are the particle numbers at site $i,j$, which are the sites linked by the edge of the physical lattice intersected with bond $a$, as noted in Fig.5 b). The analogous expressions also hold for the green and blue lattices. At filling $2/3$, the cluster-charging constraints impose “two-dimers per site” conditions on the colores honeycomb lattice with double-occupation of dimers allowed (illustrated in Fig.5 b) for the red lattice). The corresponding dimer model of each colored honeycomb lattice is the so-called “double-dimer” model, because the configuration can be thought of as overlaying two dimer coverings with one dimer touching each vertex [kenyon2014conformal].
@@
Now we introduce the height fields. In the dimer model, height fields are integer-valued fields defined at the centers of the plaquettes to encode the dimer constraints [blote1982roughening, nienhuis1984triangular]. Since the double-dimer configuration can be decomposed into two single-dimer ones, Ref. [WilkinsPhysRevB.102.174431] introduces two height fields for each of the replicas. Here, for the cluster-charging model, only the sum of the two heights is relevant since no physical operators can distinguish the two replicas. We will call this summation “the height field” in the following. The height fields of the red double-dimer model are defined at the centers of the hexagonal plaquettes, with the following rules (Fig.5 c)): when going around the vertex emitting “Y” shaped edges clockwise, if no dimer/one dimer/two dimers is crossed, height is changed by 2-2/+1+1/+4+4. The height assignment is the same for going around another type of sublattice counterclockwise. Under such rules, the action of the $\mathcal{O}_{2}$ operator is to shift the height field of one color by 3-3 and the height field of another color by +3+3 (Fig.5 d)). Now the total $\tilde{N}\equiv N_{A}-N_{B}$ conservation proved in sectionIII maps to the global conservation of the total height fields $h_{r}+h_{g}+h_{b}$.
@@
where the $\triangleright/\triangleleft$ is the shorthand notation for summing the three height fields surrounding the site of the corresponding particle (see Fig. 5 e)), where the color labels are implicit. The constant $\frac{1}{3}$ takes care of the filling fraction. The representation of the particle number using the fields on dual lattice can also be viewed as a plaquette duality, utilized for studying various spin models with ring exchange interactions [Paramekanti2002Phys.Rev.B, Myerson-Jain2022Phys.Rev.Lett.a]. The additional link to the height fields is originated from the cluster-charging constraints. Now, the on-site Hubbard term $H_{0}$ can be written as,
@@ IV.2 Coarse-grained height fields and their properties under lattice symmetries
First, the compactification radius of $h^{r,g,b}$ is $3$ for our double-dimer model, since the uniform shift by three gives the same dimer configuration (same as the dimer model on honeycomb lattice [Fradkin2004Phys.Rev.Bb]). Therefore, the physical operators can only be $\partial h^{r,g,b}$ (and higher order derivatives), the vertex operators $\mathcal{V}_{\vec{m}}(x)=e^{i\frac{2\pi}{3}\vec{m}\cdot\vec{h}(x)}$, where $\vec{m}=(m^{r},m^{g},m^{b})$ is a vector of integers and $\vec{h}=(h^{r},h^{g},h^{b})$, and their combinations.
The transformation of the coarse-grained fields under lattice symmetries can be derived from the “flat” configurations of $h^{r,g,b}$’s. These are columnar orders of the double-dimer model. However, not all columnar order combinations are physical, since not all double-dimer configurations can be mapped to physical ones. There are nine physical columnar orders in total, related to the configuration in Fig.1 b) by lattice translations and rotations. The symmetry transformation of the coarse-grained height fields need to account for not only the symmetry properties of the microscopic heights, but also the transformation of the average height, analogous to the discussion of the dimer model [Fradkin2004Phys.Rev.Bb, fradkin2013field]. The detail is discussed in the Appendix Sec.A. We will quote the results below. Under lattice symmetry,
@@ IV.3 Effective field theory description
$\mathcal{S}_{sg}[\vec{\eta}]=\int\frac{1}{4\tilde{t}^{\prime 2}}(\partial_{\tau}\vec{\eta})^{2}+\tilde{v}^{\prime}(\partial_{\mu}\vec{\eta})^{2}-\tilde{y}^{\prime}\sum_{i}\cos(4\pi\vec{\alpha}_{i}\cdot\vec{\eta}).$ (21)
@@
Now let us focus on $\mathcal{L}_{1}$, with $\tilde{y}_{0}\rightarrow\tilde{y}_{0}+\tilde{y}$. We note that $\mathcal{L}_{1}$ resembles the effective action for quantum interface with continuous symmetry [fradkin1983roughening]. The difference is that the height field in $\mathcal{L}_{1}$ is compact, $h_{0}\sim h_{0}+3$, while in the quantum interface, the height is not compact.
[…]
$\mathcal{L}_{1,n_{0}}[\theta_{0},h_{0}]\approx i\theta_{0}\partial_{\tau}h_{0}+\tilde{t}(\partial_{\mu}\theta_{0})^{2}+\tilde{v}_{0}(\partial_{\mu}h_{0})^{2}+2\pi^{2}\tilde{y}_{0}(\delta h_{0})^{2},$ (22)
where $\delta h_{0}=h_{0}-n_{0}$ and $n_{0}$ is an integer. The path integral takes the sum over all the minima, $\mathcal{Z}_{1}\sim\int Dh_{0}D\theta_{0}\sum_{n_{0}}e^{-\int\mathcal{L}_{1,n_{0}}[\theta_{0},h_{0}]}$. Ref. [fradkin1983roughening] essentially integrated out $\theta_{0}$ and focused on the physics of $\delta h_{0}$. However, the action involving $\mathcal{L}_{1}$, for each $n_{0}$, can be viewed as the hydrodynamic action of a bosonic superfluid in the long-wavelength limit by mapping $\theta_{0}$ to the phase angle of the condensate [popov1972hydrodynamic, Popov_1988]. Hence, we identified the SN phase. The ordering of $\vec{\eta}$ that is guaranteed in the ground state from analyzing $\mathcal{S}_{sg}[\vec{\eta}]$ (Eq.21), irrespective of spontaneous symmetry breaking or lack thereof in the theory of $\theta_{0}$, accounts for the “nematic” part of the phase. The ordering of $\theta_{0}$ accounts for the “ super” part in analogy to superfluid, although this is a phase conjugate to the total sublattice polarization rather than total particle number, hence SN is not a superfluid. The $\theta_{0}$-disordered phase will be analogous to the Mott insulating state in the Bose-Hubbard model [Fisher1989Phys.Rev.B], which is a $C_{3}$-symmetry-breaking, columnar-ordered phase in our case. Hence, we established the two possible ground state phases in the phase diagram sketched in Fig.1 e).
Specifically, the two distinct ground states are:
  • $\langle e^{i\frac{2\pi}{3}h_{0}}\rangle\neq 0,~\langle e^{i\frac{4\pi}{3}\vec{\alpha}_{i}\cdot\vec{\eta}}\rangle\neq 0$: the gapped columnar ordered phase.
  • $\langle e^{i\theta_{0}}\rangle\neq 0,~\langle e^{i\frac{4\pi}{3}\vec{\alpha}_{i}\cdot\vec{\eta}}\rangle\neq 0$: the gapless supernematic phase.
We discuss the properties of each phase in the following sections.
A complete theory of the quantum phase transition between the two phases requires addressing several subtleties, which are beyond the scope of the present paper. The central issue is that the vertex operators $e^{i\alpha h_{0}}$ and $e^{ip\theta_{0}}$ do not commute. As a result, the symmetry of $\mathcal{L}_{1}$ takes the nontrivial form $U(1)\rtimes Z_{3}$, with group elements $e^{i\alpha h_{0}}$ ($\alpha\in[0,2\pi)$) and $e^{ip\theta_{0}}$ ($p\in{0,1,2}$), respectively. The impossibility of simultaneously ordering these two vertex operators introduces an additional subtlety that was previously unnoticed. Formally integrating out the $h_{0}$ field in Eq.22 as one would in the hydrodynamic theory of superfluid yields the phase-only action:
@@
where $\tilde{\mathbf{y}}$ is a monotonic function of $\tilde{y}_{0}+\tilde{y}$. The Euclidean action $\mathcal{S}_{1}[\theta_{0}]$ is analogous to the 3D classical XY model with $\theta_{0}$ as the spin azimuthal angle. If the vortices of $\theta_{0}$ can be taken into account the same way as in the XY model, there would be a phase transtion between a superfluid phase and a gapped Mott insulating phases by tuning the effective stiffness $K_{\text{eff}}=\sqrt{\tilde{t}/\tilde{\mathbf{y}}}$ across a critical value $K_{\text{eff}}^{c}$[WilsonFisher]. We leave the full theory of the phase transition upon introducing vortices of $\theta_{0}$ to a future work.
@@ V Columnar ordered phase
In the classical limit of $\tilde{t}=0$, we reproduce the trimer liquid[zhang2024bionicfractionalizationtrimermodel], a classical ensemble of extensively-many degenerate configurations (1 e)). In general, small quantum fluctuations around extensive degeneracy can result in quantum order-by-disorder, where a charge ordering state is favored due to the virtual exchange processes. To determine the symmetry-breaking pattern of such a charge-ordered state and its thermal behavior, we use the effective field theory and the symmetry representations of the fields.
@@ V.2 Finite temperature
(a) Upon increasing temperature, $h_{0}$ melts first, via a three-state Potts transition. The intermediate phase has $\vec{\eta}$ ordering. Since $h_{0}$ is disordered, the order parameters are the lowest order vertex operators involving only $\vec{\eta}$-field: $\langle\mathcal{V}_{(1,-1,0)}\rangle$, $\langle\mathcal{V}_{(0,1,-1)}\rangle$ and $\langle\mathcal{V}_{(-1,0,1)}\rangle$. They all take the same value of $1,e^{i\frac{2\pi}{3}},\text{or}e^{-i\frac{2\pi}{3}}$ from Eq.24. From Eq.18, the ordering pattern breaks $\mathcal{C}_{3}$ rotation while preserving translation and inversion. Hence, this intermediate phase is three-fold degenerate, which we call “nematic”. In the disordered phase of $h_{0}$, the bionic monomers can acquire non-zero fugacity. The scaling dimension of these monomer operators can be obtained from the same method as in the classical trimer liquid [zhang2024bionicfractionalizationtrimermodel] and $\Delta_{m_{\eta}}=\frac{3K^{\prime}}{4\pi}$. Hence when $K^{\prime}=K_{c}^{{}^{\prime}}=\frac{8\pi}{3}$, the monomers start to become relevant. However, since $\Delta_{(2,-1,-1)}(K_{c}^{{}^{\prime}})=\frac{3}{2}<2$, the thermal phase transition is dominant by the cosine terms, where the value of the critical $K^{\prime}$ is nonuniversal, and expected to be three-state Potts like [PhysRevB.16.1217].
@@ VI.1 Zero temperature
where we ignore the dependence on the gapped $\eta$-fields and a spatially oscillating term (see SM A.3 for the derivation). Following the effective field theory of $\theta_{0}$ (Eq.23) and the standard derivation following the hydrodynamic theory of bosonic superfluid [Popov_1988] (see SM B), the dynamical structure factor $\tilde{S}(q,\omega)\sim|q|\delta(\omega-v|q|)$, where $\tilde{S}(q,\omega)=\int dte^{i\omega t}\langle\tilde{n}(q,t)\tilde{n}(-q,0)\rangle$, and $\tilde{n}(q)$ is the Fourier transform of $n_{A}(x_{A})-n_{B}(x_{B})$. Therefore, the SN is compressible with respect to the sublattice particle number $\tilde{N}$, following the compressibility sum-rule [10.1093/acprof:oso/9780198758884.003.0007] $\tilde{\kappa}=2\lim_{q\rightarrow 0}\int_{0}^{\infty}\frac{\tilde{S}(q,\omega)}{\omega}d\omega$. Hence, from the $\tilde{N}$ perspective, SN is compressible although it is an incompressible state for the total particle number $N=N_{A}+N_{B}$.